OpenStudy (anonymous):

solve the following system of equation x+3y=2, x=6-3y

OpenStudy (anonymous):

im trying to figure this out ok

OpenStudy (anonymous):

i know that x=3

OpenStudy (anonymous):

and y = 6

OpenStudy (kc_kennylau):

No, 3+3*6 is not 2

OpenStudy (kc_kennylau):

There is no solution: the two lines are parallel to each other

OpenStudy (mrnood):

rewrite as: x+3y=2, x=6-3y y=2/3-x/3 and y=2-x/3 The slope is the same (because the value before the x is -1/3 in both equations) Therefore the lines are parallel

OpenStudy (kc_kennylau):

Or substitute the second equation into the first: 6-3y+3y=2 6=2 Which makes it invalid Which makes there be no solution

OpenStudy (anonymous):

6=2

OpenStudy (mrnood):

@cece1902 BUT 6 is NOT equal to 2 - so see @kc-kennylau 's answer above...